What is the critical path for the network below?

End
Answer #1

This project has three network paths

Path 1= Start -B-G-I-End = 4+8+ 6 = 18

Path 2= start-A-C-E-H-End = 6+4+6+5 = 21

Path 3 = Start-A-D-F-H-End = 6+5+6+5 = 22

The critical path in this network is start - A-D-F-H-End with the path value of 22.